OpenStudy (anonymous):

Graphical Damped Osccilation System by using Graphical User Interface (GUI) on MATLAB, check it out

OpenStudy (anonymous):

function varargout = osilasi(varargin) % OSILASI MATLAB code for osilasi.fig % OSILASI, by itself, creates a new OSILASI or raises the existing % singleton*. % % H = OSILASI returns the handle to a new OSILASI or the handle to % the existing singleton*. % % OSILASI('CALLBACK',hObject,eventData,handles,...) calls the local % function named CALLBACK in OSILASI.M with the given input arguments. % % OSILASI('Property','Value',...) creates a new OSILASI or raises the % existing singleton*. Starting from the left, property value pairs are % applied to the GUI before osilasi_OpeningFcn gets called. An % unrecognized property name or invalid value makes property application % stop. All inputs are passed to osilasi_OpeningFcn via varargin. % % *See GUI Options on GUIDE's Tools menu. Choose "GUI allows only one % instance to run (singleton)". % % See also: GUIDE, GUIDATA, GUIHANDLES % Edit the above text to modify the response to help osilasi % Last Modified by GUIDE v2.5 08-Jan-2013 11:29:53 % Begin initialization code - DO NOT EDIT gui_Singleton = 1; gui_State = struct('gui_Name', mfilename, ... 'gui_Singleton', gui_Singleton, ... 'gui_OpeningFcn', @osilasi_OpeningFcn, ... 'gui_OutputFcn', @osilasi_OutputFcn, ... 'gui_LayoutFcn', [] , ... 'gui_Callback', []); if nargin && ischar(varargin{1}) gui_State.gui_Callback = str2func(varargin{1}); end if nargout [varargout{1:nargout}] = gui_mainfcn(gui_State, varargin{:}); else gui_mainfcn(gui_State, varargin{:}); end % End initialization code - DO NOT EDIT % --- Executes just before osilasi is made visible. function osilasi_OpeningFcn(hObject, eventdata, handles, varargin) % This function has no output args, see OutputFcn. % hObject handle to figure % eventdata reserved - to be defined in a future version of MATLAB % handles structure with handles and user data (see GUIDATA) % varargin command line arguments to osilasi (see VARARGIN) % Choose default command line output for osilasi handles.output = hObject; % Update handles structure guidata(hObject, handles); % UIWAIT makes osilasi wait for user response (see UIRESUME) % uiwait(handles.figure1); % --- Outputs from this function are returned to the command line. function varargout = osilasi_OutputFcn(hObject, eventdata, handles) % varargout cell array for returning output args (see VARARGOUT); % hObject handle to figure % eventdata reserved - to be defined in a future version of MATLAB % handles structure with handles and user data (see GUIDATA) % Get default command line output from handles structure varargout{1} = handles.output; function m_Callback(hObject, eventdata, handles) % hObject handle to m (see GCBO) % eventdata reserved - to be defined in a future version of MATLAB % handles structure with handles and user data (see GUIDATA) % Hints: get(hObject,'String') returns contents of m as text % str2double(get(hObject,'String')) returns contents of m as a double % --- Executes during object creation, after setting all properties. function m_CreateFcn(hObject, eventdata, handles) % hObject handle to m (see GCBO) % eventdata reserved - to be defined in a future version of MATLAB % handles empty - handles not created until after all CreateFcns called % Hint: edit controls usually have a white background on Windows. % See ISPC and COMPUTER. if ispc set(hObject,'BackgroundColor','white'); else set(hObject,'BackgroundColor',get(0,'defaultUicontrolBackgroundColor')); end function c_Callback(hObject, eventdata, handles) % hObject handle to c (see GCBO) % eventdata reserved - to be defined in a future version of MATLAB % handles structure with handles and user data (see GUIDATA) % Hints: get(hObject,'String') returns contents of c as text % str2double(get(hObject,'String')) returns contents of c as a double % --- Executes during object creation, after setting all properties. function c_CreateFcn(hObject, eventdata, handles) % hObject handle to c (see GCBO) % eventdata reserved - to be defined in a future version of MATLAB % handles empty - handles not created until after all CreateFcns called % Hint: edit controls usually have a white background on Windows. % See ISPC and COMPUTER. if ispc set(hObject,'BackgroundColor','white'); else set(hObject,'BackgroundColor',get(0,'defaultUicontrolBackgroundColor')); end function k_Callback(hObject, eventdata, handles) % hObject handle to k (see GCBO) % eventdata reserved - to be defined in a future version of MATLAB % handles structure with handles and user data (see GUIDATA) % Hints: get(hObject,'String') returns contents of k as text % str2double(get(hObject,'String')) returns contents of k as a double % --- Executes during object creation, after setting all properties. function k_CreateFcn(hObject, eventdata, handles) % hObject handle to k (see GCBO) % eventdata reserved - to be defined in a future version of MATLAB % handles empty - handles not created until after all CreateFcns called % Hint: edit controls usually have a white background on Windows. % See ISPC and COMPUTER. if ispc set(hObject,'BackgroundColor','white'); else set(hObject,'BackgroundColor',get(0,'defaultUicontrolBackgroundColor')); end function u_Callback(hObject, eventdata, handles) % hObject handle to u (see GCBO) % eventdata reserved - to be defined in a future version of MATLAB % handles structure with handles and user data (see GUIDATA) % Hints: get(hObject,'String') returns contents of u as text % str2double(get(hObject,'String')) returns contents of u as a double % --- Executes during object creation, after setting all properties. function u_CreateFcn(hObject, eventdata, handles) % hObject handle to u (see GCBO) % eventdata reserved - to be defined in a future version of MATLAB % handles empty - handles not created until after all CreateFcns called % Hint: edit controls usually have a white background on Windows. % See ISPC and COMPUTER. if ispc set(hObject,'BackgroundColor','white'); else set(hObject,'BackgroundColor',get(0,'defaultUicontrolBackgroundColor')); end function u0_Callback(hObject, eventdata, handles) % hObject handle to u0 (see GCBO) % eventdata reserved - to be defined in a future version of MATLAB % handles structure with handles and user data (see GUIDATA) % Hints: get(hObject,'String') returns contents of u0 as text % str2double(get(hObject,'String')) returns contents of u0 as a double % --- Executes during object creation, after setting all properties. function u0_CreateFcn(hObject, eventdata, handles) % hObject handle to u0 (see GCBO) % eventdata reserved - to be defined in a future version of MATLAB % handles empty - handles not created until after all CreateFcns called % Hint: edit controls usually have a white background on Windows. % See ISPC and COMPUTER. if ispc set(hObject,'BackgroundColor','white'); else set(hObject,'BackgroundColor',get(0,'defaultUicontrolBackgroundColor')); end % --- Executes on button press in Calculate. function calculate_Callback(hObject, eventdata, handles) % hObject handle to Calculate (see GCBO) % eventdata reserved - to be defined in a future version of MATLAB % handles structure with handles and user data (see GUIDATA) global m; global k; global c; global u; global u0; global wd; global neta; global x; m=str2double(get(handles.m,'string')); k=str2double(get(handles.k,'string')); c=str2double(get(handles.c,'string')); u=str2double(get(handles.u,'string')); u0=str2double(get(handles.u0,'string')); if m<=0 errordlg('Please input valid value for Mass','error','modal'); end if k<=0 errordlg('Please input valid value for Stiffness','error','modal'); end wd=(k/m)^0.5; neta=c/(2.*(m.*k)^0.5); wdd=sprintf('%0.2f',wd); netaa=sprintf('%0.2f',neta); set(handles.wd,'string',wdd); set(handles.neta,'string',netaa); t=[0:0.01:100]; if neta==1 set(handles.dis,'string','CRITICAL DAMPING'); a=u-1; b=u0+neta.*wd.*a-neta.*wd; x=1+a.*(2.72.^(-neta.*wd.*t))+b.*t.*(2.72.^(-neta.*wd.*t)); axes(handles.axes); plot(t,x,'r'); %for i = 1:1000 %m = x(1,i); %imx=[2 3 3 2]; %imy=[2+m 2+m 3+m 3+m]; %axes(handles.image); %plot(imx,imy); %pause(.5); %end elseif neta>1 set(handles.dis,'string','OVERDAMPING'); a=(1+wd.*(neta+(neta.^2-1).^0.5).*(u-1)+u0)/(2.*wd.*(neta.^2-1).^0.5); b=(u0+wd.*(neta-(neta.^2-1).^0.5).*(u-1))/(2.*wd.*(neta.^2-1).^0.5); x=1+a.*2.72.^(-wd.*(neta-(neta.^2-1).^0.5).*t)+b.*2.72.^(-wd.*(neta+(neta.^2-1).^0.5).*t); axes(handles.axes); plot(t,x,'r'); elseif neta<1 set(handles.dis,'string','UNDERDAMPING'); a=u-1; b=(u0+neta.*wd.*(u-1))/(wd.*(1-neta.^2).^0.5); x=1+2.72.^(-neta.*wd.*t).*(a.*cos((wd.*(1-neta.^2).^0.5).*t)+b.*sin((wd.*(1-neta.^2).^0.5).*t)); axes(handles.axes); plot(t,x,'r'); end %clear all; % --- Executes on button press in exit. function exit_Callback(hObject, eventdata, handles) % hObject handle to exit (see GCBO) % eventdata reserved - to be defined in a future version of MATLAB % handles structure with handles and user data (see GUIDATA) close;